# Keeping Score: A Quantitative Analysis of How the CHI Community Appreciates Its Milestones

### v2

@@ 4.1. Milestones cited by the CHI Community (RQ1 and RQ2)
The CHI community’s 30 most cited milestones with year of publication and citation count are listed in Table 1. Half of the 30 milestones most cited at CHI are published at the CHI Conference, the other half in other conferences, journals, or books. The distribution of the CHI community’s citations to milestones is long-tailed. Very few milestones received many citations, and there is a stark difference between CHI’s top performing milestones and the rest of the milestones (cf. Table 1). This demonstrates how much the CHI community’s attention rests on a few top-performing papers. Table 1 also highlights how relatively few citations the milestones are receiving from CHI authors. For instance, the thirtieth paper in the top-30 list has received only 75 citations from the CHI community since its publication in 2006. Expanding this list to the top-300 most cited papers, the papers in 300th place each received only 34 citations from the CHI community. The following section provides a description of CHI’s top-30 most cited milestones.
@@ 4.1.1. Milestone description (RQ1)
Figure 2. Relative number of citations to CHI’s overall top-100 most cited milestones over time. A clustering analysis reveals three types of milestones in CHI: one “super milestone” continues to grow in popularity (red), some past super milestones are fading from the CHI community’s memory (light blue), and some contemporary milestones are receiving continued attention, although at lower levels (blue).
[…]
Line plot of citations from CHI articles to the top-100 milestones over time, graphically showcasing three different types of milestones: contemporary milestones, super milestones, and fading super milestones.
@@
[…]
@@
[…]
Description and contributions of each of the 30 milestones most-cited by the CHI community. No. Year Milestone Description Contribution 1. (2006)   Braun and Clarke (Braun and Clarke, 2006) Thematic Analysis, a method for analyzing qualitative data method 2. (1988)   Hart and Staveland (Hart and Staveland, 1988) NASA Task Load Index (NASA-TLX), a tool for measuring perceived workload method/tool 3. (2011)   Wobbrock et al. (Wobbrock et al., 2011) Aligned Rank Transform (ART) for non-parametric factorial analyses through ANOVA procedures method 4. (2014)   Zimmerman and Forlizzi (Zimmerman and Forlizzi, 2014) Research through Design as a way to generate knowledge by reflecting on design practice method 5. (2010)   Bardzell (Bardzell, 2010) Feminist HCI, a critical agenda that considers gender, power, and identity in design framework/agenda 6. (2012)   Braun and Clarke (Braun and Clarke, 2012) Thematic Analysis, a method for analyzing qualitative data method/guidelines 7. (2019)   Braun and Clarke (Braun and Clarke, 2019) Reflexive Thematic Analysis, a clarification of Braun and Clarke’s initial method method/guidelines 8. (2006)   Hart (Hart, 2006) A reflection on the NASA-TLX’s two decades of impact method/tool 9. (2019)   McDonald et al. (McDonald et al., 2019) Guidelines on reliability and inter-rater reliability in qualitative research guidelines 10. (2010)   Li et al. (Li et al., 2010) A stage-based model of personal informatics systems model 11. (2003)   Hutchinson et al. (Hutchinson et al., 2003) Technology Probes, an expansion of Cultural Probes method 12. (2005)   Sengers et al. (Sengers et al., 2005) Reflective Design, an approach to question underlying assumptions in technology framework 13. (1996)   Brooke (Brooke, 1996) System Usability Scale (SUS), a tool for assessing usability method/tool 14. (2003)   Gaver et al. (Gaver et al., 2003) Ambiguity as a Resource for Design framework/concept 15. (2017)   Schlesinger et al. (Schlesinger et al., 2017) Intersectional HCI, a critical perspective advocating for designs that address the complexities of gender, race, and class framework 16. (1997)   Ishii and Ullmer (Ishii and Ullmer, 1997) Tangible Bits, a vision for seamless interfaces that blend the digital and physical worlds, laying the groundwork for tangible interaction concept 17. (2010)   DiSalvo et al. (DiSalvo et al., 2010) Sustainable HCI, urging the field to consider environmental impacts framework 18. (2010)   Irani et al. (Irani et al., 2010) Postcolonial Computing, a framework for examining design and development through postcolonial theory framework 19. (2014)   Rooksby et al. (Rooksby et al., 2014) The concept of “lived informatics”, examining how people integrate personal tracking into their daily lives model 20. (2019)   Amershi et al. (Amershi et al., 2019) Guidelines for Human-AI Interaction, offering best practices for designing user-centric AI systems guidelines 21. (2013)   Dunne and Raby (Dunne and Raby, 2013) Speculative Design, a method using fictional scenarios to explore alternative futures framework/concept 22. (1999)   Gaver et al. (Gaver et al., 1999) Cultural Probes, a method using evocative artifacts to engage people in the design process method 23. (2015)   Epstein et al. (Epstein et al., 2015) Lived Informatics Model, a development of Rooksby et al.’s method model 24. (2016)   Azmandian et al. (Azmandian et al., 2016) Haptic Retargeting, a technique for creating more immersive interactions in VR experiences method 25. (2009)   Wobbrock et al. (Wobbrock et al., 2009) User-defined gestures, underscoring the importance of user preferences in surface computing interactions method 26. (2013)   Follmer et al. (Follmer et al., 2013) inFORM, a shape display that allows for dynamic physical affordances through actuation, demonstrating the potential of tangible interfaces in HCI method 27. (1993)   Kennedy et al. (Kennedy et al., 1993) Simulator Sickness Questionnaire, a method for quantifying simulator sickness method/tool 28. (2011)   Bostock et al. (Bostock et al., 2011) D3 (Data-Driven Documents), a software library for creating interactive web visualizations tool 29. (2003)   MacKenzie and Soukoreff (MacKenzie and Soukoreff, 2003) Standardized phrase sets for evaluating text entry techniques as benchmark for assessing typing performance method/tool 30. (2006)   Charmaz (Charmaz, 2006) A practical guide to Constructing Grounded Theory method
@@
Next, we turn our attention to analyzing the different contribution types the milestones make. This is crucial for understanding what type of work is being appreciated in the long run.
@@ 4.1.2. Milestone contributions (RQ2)
[…]
Comparison of milestone contribution types with existing contribution types in HCI. Contribution Description with examples Wobbrock and Kientz ((2016))* ACM SIGCHI ((2024)$\dagger$ Method A systematic approach or technique used for conducting research or analysis in HCI.Examples: qualitative methods (e.g., thematic analysis), quantitative methods (e.g., aligned rank transform) Methodological Methodology Tool A practical instrument or software designed to perform specific tasks or measurements in HCI studies. Examples: survey tools (e.g., NASA-TLX, SUS), software libraries (e.g., D3) Artifact Artifacts Framework A comprehensive structure outlining key concepts in HCI which organizes knowledge and provides a structured approach for HCI research. Examples: Feminist HCI, Reflective Design Theoretical Theory Agenda A call for future research or practice directions in HCI.Example: Feminist HCI agenda Opinion Argument Guidelines Prescriptive recommendations or best practices for designing or evaluating human-computer interaction. Example: Guidelines for inter-rater reliability Methodological Methodology Model An abstract representation detailing the components and processes of a system or phenomenon in HCI. Example: Model of personal informatics Theoretical Theory Concept A foundational idea or theory that introduces new perspectives or approaches to HCI design and research.Example: Ambiguity Theoretical Theory
* Research contribution types in HCI (Wobbrock and Kientz, 2016): Empirical, Artifact, Methodological, Theoretical, Dataset, Survey, Opinion $\dagger$ Contributions to CHI (ACM SIGCHI, 2024): 1) Development or Refinement of Interface Artifacts or Techniques, 2) Understanding Users, 3) Systems, Tools, Architectures, and Infrastructure, 4) Methodology, 5) Theory, 6) Innovation, Creativity, and Vision, 7) Argument, 8) Validation and Replication
@@ 4.1.3. Case study: The rise of CHI’s super milestone
Braun and Clarke’s Thematic Analysis (Braun and Clarke, 2006) is a unique “super” milestone (cf. Figure 5 and Figure 6). What makes this milestone stand out is not only its absolute citation count, but also its speed of adoption in the CHI community. Braun and Clarke’s method for qualitative analysis of data was first published in the year (2006) and was immediately useful to researchers in HCI, resulting in rapid adoption in the CHI community (cf. Figure 2 and (Bowman et al., 2023)). In fact, Thematic Analysis occupies not one but three of CHI’s 30 most cited milestones (see Table 1). Together, these milestone papers have amassed 1024 citations from CHI papers. While this number pales in comparison to the total number of citations that these milestone papers have received from other sources, it does demonstrate the popularity of Thematic Analysis in the CHI community. Braun and Clarke’s milestones have received almost four times more citations than the second most cited milestone at CHI (see Table 1).
@@ 4.2. How has the CHI Community’s Attention to its Milestones evolved over time? (RQ3)
[…]
Line plot of variance of derivative and total variation of the CHI forgetting curves.
@@
[…]
Cumulative citations to the prior 300 milestones from selected CHI proceedings years.
In the 2024 Proceedings of CHI, a significant portion of the community’s attention is directed toward more recent contributions. To understand this trend, we calculated the cumulative percentage of citations from selected CHI proceedings to the 300 most cited milestones published in years prior (see Figure 4). The results highlight the concentration of citations to recent milestones. Notably, 15.4% of the citations point to works published in 2023. This cumulative percentage increases to 27.5% when including works from 2022 and 2023, and further rises to 60.1% when considering all works from 2018 onwards. By 2010, the cumulative percentage reaches 79.6%, and by 2005, it climbs to 89.9%. The bulk of the CHI community’s attention rests on papers published in a time window of between 5 and 15 years. This indicates a strong emphasis on citing more recent milestones within the CHI community.
@@
[…]
A plot depicting how the CHI community’s citations to the 300 most cited prior milestones have developed over time. Looking at the panels, we can note that milestones have lost influence over time at CHI, relative to all other citations in a given proceedings year.
@@
[…]
Milestone co-efficient plotted for the 3,000 papers most cited at CHI.
We apply the milestone coefficient (MC) to the top-30 milestones (as listed in Table 1). The MC in Table 1 is negative only for one milestone (D3 (Bostock et al., 2011)) which has seen a decline in citations from the CHI community in recent years. All other milestones in the top-30 list are still trending toward growth in citations. To extend this analysis and derive insights on the CHI as a whole, we apply the milestone coefficient to the 3,000 papers most cited by CHI authors. For these papers, the milestone coefficient ranges from -0.0102 to 0.0768. Figure 6 plots the MC for the top 3,000 most cited papers. Each bubble represents one paper in its year of publication. The size of the bubble represents the total number of citations to this milestone. The exceptional super milestone is clearly visible in the year 2006, achieving an MC of 0.0768 (the highest MC in the dataset). The second most cited milestone, the NASA TLX (Hart and Staveland, 1988), is visible in 1988. From Figure 6, it can be derived that the number of milestones has grown over the years. This is unsurprising, given the growth of the CHI proceedings and the growth in citations (Oppenlaender, 2025). However, the figure also demonstrates that many recent milestones struggle to receive the same attention from the CHI community as the past super milestones. Some of the recent milestones even trend towards a decline in citations, and the number of milestones that are being ‘forgotten’ has increased in the last two decades.

### v3

@@ 4.1. Milestones Cited by the CHI Community (RQ1 and RQ2)
The CHI community’s 30 most cited milestones with year of publication and citation count are listed in Table 1. Half of the 30 milestones most cited at CHI are published at the CHI Conference, the other half in other conferences, journals, or books. The distribution of the CHI community’s citations to milestones is long-tailed. Very few milestones received many citations, and there is a stark difference between CHI’s top performing milestones and the rest of the milestones (cf. Table 1). This demonstrates how much the CHI community’s attention rests on a few top-performing papers. Table 1 also highlights how relatively few citations the CHI community’s milestones are receiving from CHI authors. For instance, the thirtieth paper in the top-30 list has received only 75 citations from the CHI community since its publication in 2006. Expanding this list to the top-300 most cited papers, the papers in 300th place each received only 34 citations from the CHI community (see the Supplemental Material). The following section provides a description of CHI’s top-30 most cited milestones.
@@ 4.1.1. Milestone description (RQ1)
1. 2006   Braun and Clarke (Braun and Clarke, 2006) Using thematic analysis in psychology 770 journal 7.68
2. 1988   Hart and Staveland (Hart and Staveland, 1988) Development of NASA-TLX (task load index): results of empirical and theoretical research 262 book 2.08
3. 2011   Wobbrock et al. (Wobbrock et al., 2011) The aligned rank transform for nonparametric factorial analyses using only ANOVA procedures 179 CHI 1.66
4. 2007   Zimmerman and Forlizzi (Zimmerman and Forlizzi, 2014) Research through design as a method for interaction design research in HCI 162 book 1.38
5. 2010   Bardzell (Bardzell, 2010) Feminist HCI: taking stock and outlining an agenda for design 141 CHI 1.19
6. 2012   Braun and Clarke (Braun and Clarke, 2012) Thematic analysis 129 book 1.85
7. 2019   Braun and Clarke (Braun and Clarke, 2019) Reflecting on reflexive thematic analysis 125 book 2.22
8. 2006   Hart (Hart, 2006) NASA-task load index (NASA-TLX); 20 years later 113 journal 1.35
9. 2019   McDonald et al. (McDonald et al., 2019) Reliability and inter-rater reliability in qualitative research: norms and guidelines for CSCW and HCI practice 113 CHI 1.68
10. 2010   Li et al. (Li et al., 2010) A stage-based model of personal informatics systems 107 CHI 1.01
11. 2003   Hutchinson et al. (Hutchinson et al., 2003) Technology probes: inspiring design for and with families 106 CHI 0.98
12. 2005   Sengers et al. (Sengers et al., 2005) Reflective design 97 CC 0.82
13. 1996   Brooke (Brooke, 1996) SUS: a ‘quick and dirty’ usability scale 94 book 0.98
14. 2003   Gaver et al. (Gaver et al., 2003) Ambiguity as a resource for design 92 CHI 0.76
15. 2017   Schlesinger et al. (Schlesinger et al., 2017) Intersectional HCI: engaging identity through gender, race, and class 86 CHI 0.90
16. 1997   Ishii and Ullmer (Ishii and Ullmer, 1997) Tangible bits: towards seamless interfaces between people, bits and atoms 84 CHI 0.77
17. 2010   DiSalvo et al. (DiSalvo et al., 2010) Mapping the landscape of sustainable HCI 84 CHI 0.85
18. 2010   Irani et al. (Irani et al., 2010) Postcolonial computing: a lens on design and development 84 CHI 0.77
19. 2014   Rooksby et al. (Rooksby et al., 2014) Personal tracking as lived informatics 84 CHI 0.59
20. 2019   Amershi et al. (Amershi et al., 2019) Guidelines for human-AI interaction 84 CHI 0.60
21. 2013   Dunne and Raby (Dunne and Raby, 2013) Speculative everything: design, fiction, and social dreaming 83 book 0.96
22. 1999   Gaver et al. (Gaver et al., 1999) Design: Cultural probes 80 journal 0.73
23. 2015   Epstein et al. (Epstein et al., 2015) A lived informatics model of personal informatics 78 UbiComp 0.78
24. 2016   Azmandian et al. (Azmandian et al., 2016) Haptic retargeting: dynamic repurposing of passive haptics for enhanced virtual reality experiences 78 CHI 0.65
25. 2009   Wobbrock et al. (Wobbrock et al., 2009) User-defined gestures for surface computing 77 CHI 0.65
26. 2013   Follmer et al. (Follmer et al., 2013) inFORM: dynamic physical affordances and constraints through shape and object actuation 77 UIST 0.75
27. 1993   Kennedy et al. (Kennedy et al., 1993) Simulator sickness questionnaire: an enhanced method for quantifying simulator sickness 76 journal 0.62
28. 2011   Bostock et al. (Bostock et al., 2011) D3: Data-driven documents 76 journal -0.62
29. 2003   MacKenzie and Soukoreff (MacKenzie and Soukoreff, 2003) Phrase sets for evaluating text entry techniques 75 CHI 0.57
30. 2006   Charmaz (Charmaz, 2006) Constructing grounded theory a practical guide through qualitative analysis 75 book 0.77
@@
No. Year Milestone Description Contribution
1. (2006)   Braun and Clarke (Braun and Clarke, 2006) Thematic Analysis, a method for analyzing qualitative data method
2. (1988)   Hart and Staveland (Hart and Staveland, 1988) NASA Task Load Index (NASA-TLX), a tool for measuring perceived workload method/tool
3. (2011)   Wobbrock et al. (Wobbrock et al., 2011) Aligned Rank Transform (ART) for non-parametric factorial analyses through ANOVA procedures method
4. (2014)   Zimmerman and Forlizzi (Zimmerman and Forlizzi, 2014) Research through Design as a way to generate knowledge by reflecting on design practice method
5. (2010)   Bardzell (Bardzell, 2010) Feminist HCI, a critical agenda that considers gender, power, and identity in design framework/agenda
6. (2012)   Braun and Clarke (Braun and Clarke, 2012) Thematic Analysis, a method for analyzing qualitative data method/guidelines
7. (2019)   Braun and Clarke (Braun and Clarke, 2019) Reflexive Thematic Analysis, a clarification of Braun and Clarke’s initial method method/guidelines
8. (2006)   Hart (Hart, 2006) A reflection on the NASA-TLX’s two decades of impact method/tool
9. (2019)   McDonald et al. (McDonald et al., 2019) Guidelines on reliability and inter-rater reliability in qualitative research guidelines
10. (2010)   Li et al. (Li et al., 2010) A stage-based model of personal informatics systems model
11. (2003)   Hutchinson et al. (Hutchinson et al., 2003) Technology Probes, an expansion of Cultural Probes method
12. (2005)   Sengers et al. (Sengers et al., 2005) Reflective Design, an approach to question underlying assumptions in technology framework
13. (1996)   Brooke (Brooke, 1996) System Usability Scale (SUS), a tool for assessing usability method/tool
14. (2003)   Gaver et al. (Gaver et al., 2003) Ambiguity as a Resource for Design framework/concept
15. (2017)   Schlesinger et al. (Schlesinger et al., 2017) Intersectional HCI, a critical perspective advocating for designs that address the complexities of gender, race, and class framework
16. (1997)   Ishii and Ullmer (Ishii and Ullmer, 1997) Tangible Bits, a vision for seamless interfaces that blend the digital and physical worlds, laying the groundwork for tangible interaction concept
17. (2010)   DiSalvo et al. (DiSalvo et al., 2010) Sustainable HCI, urging the field to consider environmental impacts framework
18. (2010)   Irani et al. (Irani et al., 2010) Postcolonial Computing, a framework for examining design and development through postcolonial theory framework
19. (2014)   Rooksby et al. (Rooksby et al., 2014) The concept of “lived informatics”, examining how people integrate personal tracking into their daily lives model
20. (2019)   Amershi et al. (Amershi et al., 2019) Guidelines for Human-AI Interaction, offering best practices for designing user-centric AI systems guidelines
21. (2013)   Dunne and Raby (Dunne and Raby, 2013) Speculative Design, a method using fictional scenarios to explore alternative futures framework/concept
22. (1999)   Gaver et al. (Gaver et al., 1999) Cultural Probes, a method using evocative artifacts to engage people in the design process method
23. (2015)   Epstein et al. (Epstein et al., 2015) Lived Informatics Model, a development of Rooksby et al.’s method model
24. (2016)   Azmandian et al. (Azmandian et al., 2016) Haptic Retargeting, a technique for creating more immersive interactions in VR experiences method
25. (2009)   Wobbrock et al. (Wobbrock et al., 2009) User-defined gestures, underscoring the importance of user preferences in surface computing interactions method
26. (2013)   Follmer et al. (Follmer et al., 2013) inFORM, a shape display that allows for dynamic physical affordances through actuation, demonstrating the potential of tangible interfaces in HCI method
27. (1993)   Kennedy et al. (Kennedy et al., 1993) Simulator Sickness Questionnaire, a method for quantifying simulator sickness method/tool
28. (2011)   Bostock et al. (Bostock et al., 2011) D3 (Data-Driven Documents), a software library for creating interactive web visualizations tool
29. (2003)   MacKenzie and Soukoreff (MacKenzie and Soukoreff, 2003) Standardized phrase sets for evaluating text entry techniques as benchmark for assessing typing performance method/tool
30. (2006)   Charmaz (Charmaz, 2006) A practical guide to Constructing Grounded Theory method
@@
Next, we turn our attention to analyzing the different contribution types made by the milestones. This is crucial for understanding what type of work is being appreciated in the long run.
@@ 4.1.2. Milestone contributions (RQ2)
Contribution Description with examples Wobbrock and Kientz ((2016))* ACM SIGCHI ((2024)$\dagger$
Method A systematic approach or technique used for conducting research or analysis in HCI.Examples: qualitative methods (e.g., thematic analysis), quantitative methods (e.g., aligned rank transform) Methodological Methodology
Tool A practical instrument or software designed to perform specific tasks or measurements in HCI studies. Examples: survey tools (e.g., NASA-TLX, SUS), software libraries (e.g., D3) Artifact Artifacts
Framework A comprehensive structure outlining key concepts in HCI which organizes knowledge and provides a structured approach for HCI research. Examples: Feminist HCI, Reflective Design Theoretical Theory
Agenda A call for future research or practice directions in HCI.Example: Feminist HCI agenda Opinion Argument
Guidelines Prescriptive recommendations or best practices for designing or evaluating human-computer interaction. Example: Guidelines for inter-rater reliability Methodological Methodology
Model An abstract representation detailing the components and processes of a system or phenomenon in HCI. Example: Model of personal informatics Theoretical Theory
Concept A foundational idea or theory that introduces new perspectives or approaches to HCI design and research.Example: Ambiguity Theoretical Theory
* Research contribution types in HCI (Wobbrock and Kientz, 2016): Empirical, Artifact, Methodological, Theoretical, Dataset, Survey, Opinion
$\dagger$ Contributions to CHI (ACM SIGCHI, 2024): 1) Development or Refinement of Interface Artifacts or Techniques, 2) Understanding Users, 3) Systems, Tools, Architectures, and Infrastructure, 4) Methodology, 5) Theory, 6) Innovation, Creativity, and Vision, 7) Argument, 8) Validation and Replication
@@ 4.1.3. Case study: The rise of CHI’s super milestone
Braun and Clarke’s Thematic Analysis (Braun and Clarke, 2006) is a unique “super” milestone (cf. Figure 5 and Figure 6). What makes this milestone stand out is not only its absolute citation count, but also its speed of adoption in the CHI community. Braun and Clarke’s method for qualitative analysis was first published in the year (2006) and was immediately useful to researchers in HCI, resulting in rapid adoption in the CHI community (cf. Figure 2 and (Bowman et al., 2023)). In fact, Thematic Analysis occupies not one but three of CHI’s 30 most cited milestones (see Table 1). Together, these milestone papers have amassed 1024 citations from CHI papers. While this number pales in comparison to the total number of citations that these milestone papers have received from other sources, it does demonstrate the popularity of Thematic Analysis in the CHI community. Braun and Clarke’s milestones have received almost four times more citations than the second most cited milestone at CHI (see Table 1).
@@ 4.2. How has the CHI Community’s Attention to its Milestones evolved over time? (RQ3)
In the 2024 CHI Proceedings, a significant portion of the community’s attention is directed toward more recent contributions. To understand this trend, we calculated the cumulative percentage of citations from selected CHI proceedings to the 300 most cited milestones published in years prior (see Figure 4). The results highlight the concentration of citations to recent milestones. Notably, 15.4% of the citations point to works published in 2023. This cumulative percentage increases to 27.5% when including works from 2022 and 2023, and further rises to 60.1% when considering all works from 2018 onwards. By 2010, the cumulative percentage reaches 79.6%, and by 2005, it climbs to 89.9%. The bulk of the CHI community’s attention rests on papers published in a time window of between 5 and 15 years. This indicates a strong emphasis on citing more recent milestones within the CHI community.
@@
We apply the milestone coefficient (MC) to the top-30 milestones (as listed in Table 1). The MC in Table 1 is negative only for one milestone (D3 (Bostock et al., 2011)) which has seen a decline in citations from the CHI community in recent years. All other milestones in the top-30 list are still trending toward growth in citations. To extend this analysis and derive insights on the CHI as a whole, we apply the milestone coefficient to the 3,000 papers most cited by CHI authors. For these papers, the milestone coefficient ranges from -0.0102 to 0.0768. Figure 6 plots the MC for the top 3,000 most cited papers. Each bubble represents one paper in its year of publication. The size of the bubble represents the total number of citations to this milestone. The exceptional super milestone is clearly visible in the year 2006, achieving an MC of 0.0768 (the highest MC in the dataset). The second most cited milestone, the NASA TLX (Hart and Staveland, 1988), is visible in 1988. From Figure 6, it can be derived that the number of milestones has grown over the years. This is unsurprising, given the growth of the CHI proceedings and the growth in citations (Oppenlaender, 2025). However, the figure also demonstrates that many recent milestones struggle to receive the same attention from the CHI community as past super milestones. Some of the recent milestones even trend towards a decline in citations, and the number of milestones that are being ‘forgotten’ has increased in the last two decades.